### v1

# BE Lyncis: An Extremely Eccentric Binary with the Nearest Known Black Hole

Abstract
We report the discovery of an exceptionally eccentric binary system, BE Lyncis (BE Lyn), which hosts the nearest known black hole (BH) to Earth. Through the analysis of TESS photometry combined with an extensive set of times of maximum light spanning 39 years, we identify BE Lyn as a high-amplitude $\delta$ Scuti star in a binary with an orbital period of $\approx 15.9$ years and an extraordinary orbital eccentricity of $e=0.9989^{+0.0008}_{-0.0021}$ (>0.9968>0.9968 at 95% confidence)—the highest reliably measured for any binary system. Dynamical constraints impose an upper limit on the orbital inclination of $i\lesssim 4.0^{\circ}$, corresponding to a companion mass of $M_{2}\gtrsim 17.5~M_{\odot}$, which unequivocally favors a black hole. This system provides a unique laboratory for studying asteroseismology in strong gravitational fields, the formation of black holes via asymmetric supernovae, and the evolution of extreme binary systems. Our work demonstrates, for the first time, the successful application of the light-travel time effect in a pulsating variable to unveil a dormant black hole, establishing a novel method for BH detection in non-interacting binaries.
I Introduction
Black holes (BHs) of stellar origin represent one of the most enigmatic predictions of general relativity. While several dozen stellar-mass BHs have been identified in X-ray binaries through accretion signatures (R. W. Romani, 1998), the population of dormant (non-accreting) BHs in detached binaries remains largely unexplored. Detecting such systems is challenging, as they typically reveal themselves only through subtle gravitational interactions with luminous companions. Recent astrometric surveys, such as Gaia, have uncovered a handful of such quiescent BH binaries (K. El-Badry et al., 2023; A. Tanikawa et al., 2023; Gaia Collaboration et al., 2024), yet they remain exceedingly rare. An independent, complementary detection method—leveraging the precise clocks provided by pulsating stars—offers a promising pathway to expand this census and probe different regions of parameter space.
Pulsating $\delta$ Scuti stars are particularly well-suited for this purpose. These stars, with spectral types A–F, occupy the intersection of the main sequence and the classical Cepheid instability strip in the Hertzsprung–Russell diagram. Their oscillations, driven by the $\kappa$ mechanism operating in the partial ionization zone of helium, provide remarkably stable periodic signals with periods ranging from 15 minutes to 8 hours (T. Kallinger et al., 2008; G. Handler, 2009; D. B. Guenther et al., 2009; K. Uytterhoeven et al., 2011; D. L. Holdsworth et al., 2014; T. Steindl et al., 2022). The subclass of high-amplitude $\delta$ Scuti stars (HADS) exhibits photometric variations exceeding 0.1 mag and typically rotates slowly ($v\sin i\leq 30$ km s-1), making them ideal precision tools for detecting the minite period modulations induced by orbital motion in binary systems—the light-travel time effect (LTTE).
BE Lyncis (BE Lyn; HD 79889) is a well-studied HADS with a mean magnitude of $\langle V\rangle=8.8$ mag, spectral type A3, and sub-solar metallicity ([Fe/H] = -0.598) (C. Kim et al., 2012; Abdurro’uf et al., 2022). It was first identified as a variable star by T. Oja (1986) and later confirmed as a HADS with a pulsation period of $\sim$0.09587 days (T. Oja, 1987). Subsequent studies of its period variations revealed anomalies in OCO-C diagrams, occasionally suggestive of a parabolic trend indicative of a binary companion (e.g., E. Rodriguez et al., 1990; Y. Liu et al., 1991; E. Wunder et al., 1992; C. Boonyarak et al., 2011). Several investigations indeed proposed the existence of a low-mass companion (L. L. Kiss & K. Szatmary, 1995; A. Derekas et al., 2009; C. Boonyarak et al., 2011; J. H. Peña et al., 2015; A. Liakos & P. Niarchos, 2017), though inferred orbital parameters varied significantly due to limitations in the temporal coverage and precision of the data.
Two key developments now enable a definitive reassessment of this system. First, high-precision, continuous photometry from the Transiting Exoplanet Survey Satellite (TESS) allows for an exquisite characterization of the primary star’s pulsation modes and yields significantly more accurate times of maximum light. Second, modern computational techniques, such as robust multi-parameter global fitting via Markov Chain Monte Carlo (MCMC) methods, permit a physically consistent and precise determination of binary orbital parameters.
Here, we present a comprehensive analysis of BE Lyn, combining TESS photometry with a newly compiled, 39-year baseline of times of maximum light. Our work not only confirms its binary nature but reveals an astrophysical gem: an ultra-wide, extremely eccentric orbit harboring a massive black hole companion. This system holds multiple distinctions: (i) it possesses the highest orbital eccentricity ever reliably measured in a binary; (ii) it hosts the closest known black hole to Earth, at a distance of approximately 250 pc (C. A. L. Bailer-Jones et al., 2021); (iii) it is the only known binary comprising a pulsating star and a quiescent black hole, opening a new window for asteroseismology in strong gravity; and (iv) its extreme orbital configuration places stringent constraints on the natal kick imparted during the black hole’s formation, challenging models of asymmetric supernova explosions. The detection methodology itself—using pulsation timing in a $\delta$ Scuti star—pioneers a new, independent channel for discovering dormant black holes.
II Results and Discussion
II.1 Pulsation Properties and the Stellar Clock
Analysis of the TESS photometry (Figure A1) reveals BE Lyn as a multi-periodic pulsator. We identify 37 significant frequencies, comprising 7 independent modes and 30 harmonics or combination frequencies (Figure A2, Table A1). The two dominant frequencies, $f_{0}=10.4309\pm 0.0005$ c/d and $f_{1}=13.4226\pm 0.0005$ c/d, exhibit a period ratio of $P_{1}/P_{0}=0.777$, unambiguously identifying them as the fundamental and first-overtone radial modes, respectively. The fundamental mode amplitude dominates the light curve, exceeding that of the first overtone by a factor of $\sim 45$ and the remaining independent modes by factors >700>700 (Figure A2). This clear amplitude hierarchy renders the $f_{0}$ pulsation a remarkably clean and stable clock, enabling the application of the classical OCO-C method to detect minute orbital perturbations with negligible contamination from multi-periodic effects. The five additional independent frequencies ($f_{2}$$f_{6}$), residing in the 9179-17 c/d range with low amplitudes, are likely non-radial p-modes. This rich pulsation spectrum establishes BE Lyn as a promising target for future detailed asteroseismic modeling.
II.2 Orbital Solution: An Extraordinarily Eccentric Binary
The OCO-C analysis, incorporating a 39-year baseline of 445 times of maximum light, reveals a pronounced, long-period modulation superimposed on a secular period increase (Figure 1). Global fitting with a model combining a quadratic term (intrinsic period variation) and a light-travel time effect (LTTE) term yields the orbital solution presented in Table 1.
Figure 1: OCO-C diagram of BE Lyn (with the linear ephemeris subtracted). The top panel shows the data and the best-fit model (magenta solid line), with the quadratic component indicated by the green dashed line. The bottom panel displays the residuals. Data sources are color-coded: C. Boonyarak et al. (2011) (blue), historical literature (J. Hubscher et al., 2013; J. Hubscher, 2014; J. Hübscher, 2017; J. H. Peña et al., 2015; L. Pagel, 2021, 2022, 2023) (gray), AAVSO (red), and TESS (green).
Table 1: Pulsation and Orbital Parameters for BE Lyn.
Parameter Value
$\mathrm{BJD}_{0}$ $2449018.26853^{+0.00002}_{-0.00002}$
$P_{0}$ (d) $0.0958695052^{+0.0000000009}_{-0.0000000009}$
$\beta$ (d cycle-1) $(2.6^{+0.2}_{-0.2})\times 10^{-13}$
AA (d) $0.006^{+0.007}_{-0.003}$
$e$ $0.9989^{+0.0008}_{-0.0021}$
$P_{\mathrm{orb}}$ (d) $5811^{+29}_{-30}$
$T_{0}$ (BJD) $2447989^{+30}_{-30}$
$\omega$ () $351^{+5}_{-6}$
$(1/P_{0})(dP_{0}/dt)$ (yr-1)aaDerived from the directly fitted parameters above. $(1.02^{+0.08}_{-0.08})\times 10^{-8}$
$a\sin i$ (AU)aaDerived from the directly fitted parameters above. $1.1^{+1.3}_{-0.4}$
$f(m)$ (${M}_{\odot}$)aaDerived from the directly fitted parameters above. $0.005^{+0.046}_{-0.004}$
Note. —
The most striking result is the extraordinary orbital eccentricity of $e=0.9989^{+0.0008}_{-0.0021}$ (>0.9968>0.9968 at 95% confidence). With this value, BE Lyn claims the record for the highest reliably measured eccentricity in any binary system, surpassing previous benchmarks such as WDS 05354-3316 ($e=0.985\pm 0.002$ (A. Tokovinin et al., 2020)) and $\zeta$ Boo ($e=0.980450\pm 0.000064$ I. Waisberg et al. 2024, 2025). The orbital period is $P_{\mathrm{orb}}=5811^{+29}_{-30}$ days ($\approx 15.9$ years). The associated mass function, $f(m)=0.005^{+0.046}_{-0.004}M_{\odot}$, and small projected semi-major axis ($a\sin i=1.1^{+1.3}_{-0.4}$ AU) are indicative of a massive, unseen companion. This discovery underscores the power of long-term, precise photometric monitoring of pulsating stars to unveil extreme binary architectures.
II.3 Characterization of the Pulsating Primary
To place dynamical constraints in context, we determined the properties of the primary star through asteroseismic modeling. We computed a grid of stellar evolutionary tracks using MESA (Modules for Experiments in Stellar Astrophysics; B. Paxton et al., 2011, 2013, 2015, 2018, 2019; A. S. Jermyn et al., 2023) and corresponding adiabatic pulsation frequencies using GYRE (R. H. D. Townsend & S. A. Teitler, 2013; R. H. D. Townsend et al., 2018; J. Goldstein & R. H. D. Townsend, 2020; M. Sun et al., 2023). The optimal model, selected by matching the observed $f_{0}$, $f_{1}$, and the intrinsic period variation rate $(1/P_{0})(dP_{0}/dt)$, is summarized in Table A2 and located on the Hertzsprung-Russell diagram in Figure 2.
The best-fit model places the primary on the post-main sequence with a mass of $M_{1}=1.55M_{\odot}$, effective temperature $\log T_{\mathrm{eff}}=3.902$, and luminosity $\log(L/L_{\odot})=1.326$. The agreement between the observed $(1.02\pm 0.08)\times 10^{-8}$ yr-1 and modeled ($1.34\times 10^{-8}$ yr-1) period variation rates is noteworthy. Unlike some HADS in binaries that exhibit anomalously large period variations due to mass transfer (e.g., H.-F. Xue & J.-S. Niu, 2020; M. Lian et al., 2025), BE Lyn’s primary appears consistent with single-star evolution. This suggests the system is currently detached and has not undergone significant mass exchange.
Figure 2: Hertzsprung–Russell diagram showing evolutionary tracks (colored solid lines from the zero-age main sequence, ZAMS, to post-MS) and the best-fit seismic model for BE Lyn’s primary (red star). Black rectangles highlight zoomed-in views around the best-fit solutions. Colored dashed lines indicate tracks with a $0.01M_{\odot}$ mass step.
II.4 A Nearby Black Hole in an Extreme Orbit
Combining the orbital parameters with the primary mass ($M_{1}=1.55M_{\odot}$) introduces critical constraints from orbital stability: (i) the periastron separation must exceed the primary’s radius to avoid catastrophic collision; (ii) the primary’s Roche lobe radius must exceed the primary’s radius to avoid Roche-lobe overflow near the periastron. Applying these conditions, along with the mass function (Eq. A3), tightly constrains the system geometry (Figure 3). We derive an upper limit on the orbital inclination of $i\lesssim 4.0^{\circ}$, implying a lower limit on the companion mass of $M_{2}\gtrsim 17.5M_{\odot}$.
Figure 3: Dynamical constraints on the unseen companion. Top: Companion mass $M_{2}$ versus orbital inclination $i$, derived from the mass function (red curve). Bottom: Primary-companion separation at periastron versus $i$ (green curve). The intersection of the primary’s Roche lobe radius $R_{1,\mathrm{R}}$ (blue curve) and its physical radius $R_{1}$ (grey curve) sets the stringent upper limit $i\lesssim 4.0^{\circ}$ (dark blue dot), corresponding to the lower mass limit $M_{2}\gtrsim 17.5M_{\odot}$ (dark red dot). The gray hatched region denotes parameter space excluded.
A compact object of this mass cannot be supported by electron or neutron degeneracy pressure, ruling out a white dwarf or neutron star. The companion must therefore be a black hole. With a Gaia DR3 distance of approximately 250 pc (C. A. L. Bailer-Jones et al., 2021), this black hole becomes the nearest known to Earth, residing about half the distance of the previous record-holders Gaia BH1 (480 pc (K. El-Badry et al., 2023)) and Gaia BH3 (590 pc (Gaia Collaboration et al., 2024)). Its proximity offers unparalleled opportunities for detailed follow-up study.
The system’s most enigmatic feature—its extreme eccentricity—holds the key to its formation history. The absence of evidence for a dense stellar environment makes a dynamical scattering origin unlikely. Instead, the $e\approx 0.9989$ orbit is a natural signature of a strong natal kick imparted to the black hole during an asymmetric supernova explosion (H.-T. Janka, 2013; S. Popov et al., 2025). BE Lyn thus provides a pristine laboratory to constrain kick velocities and the physics of mass ejection in fallback supernovae, challenging models to account for the most eccentric black-hole binary known.
II.5 Broader Implications and Future Prospects
BE Lyn represents a system of singular importance across multiple astrophysical domains. First, it is the only known binary pairing a pulsating star with a quiescent black hole. This unique configuration opens the novel field of strong-gravity asteroseismology, where the pulsations of the primary could, in principle, probe the local spacetime metric perturbed by the black hole companion, especially during the brief, close periastron passages.
Second, the discovery validates the pulsation-timing method as a powerful, independent technique for detecting dormant stellar-mass black holes in wide binaries. This approach is complementary to radial-velocity and astrometric surveys, being particularly sensitive to long-period systems where other methods require extended baselines.
Key questions remain. Precise dynamical mass measurements via future astrometric missions (e.g., Gaia, Theia) or very-long-baseline interferometry could pin down $M_{2}$ and $i$. Multi-wavelength searches for accretion signatures or bow-shock emission are warranted, despite the wide separation. Theoretical work is urgently needed to simulate binary evolution pathways that yield such extreme eccentricities post-supernova.
In conclusion, BE Lyn is not merely another binary system; it is a benchmark object. It challenges our understanding of black hole formation kicks, offers a new tool for black hole discovery, and provides a unique cosmic laboratory at our galactic doorstep.
III Conclusion
We have established BE Lyncis as an extraordinary binary system. The primary is a post-main sequence high-amplitude $\delta$ Scuti star with a mass of $1.55\,M_{\odot}$. Its unseen companion orbits with an unprecedented eccentricity of $e=0.9989^{+0.0008}_{-0.0021}$ (95% confidence >>0.9968), making this the most eccentric binary known to date. This extreme orbit, coupled with dynamical constraints that demand a companion mass $M_{2}\gtrsim 17.5\,M_{\odot}$ at an orbital inclination $i\lesssim 4.0^{\circ}$, unequivocally identifies the companion as a stellar-mass black hole. With a distance of approximately 250 pc, this black hole is the closest known to Earth.
The discovery of BE Lyn carries profound implications across several astrophysical frontiers. First, its record-breaking eccentricity serves as a direct observational constraint on the strength of natal kicks imparted during asymmetric supernova explosions, challenging models of black hole formation. Second, it is the first known binary comprising a pulsating star and a quiescent black hole, creating a unique laboratory for strong-gravity asteroseismology. Third, this work validates the pulsation-timing method—using the precise clocks of $\delta$ Scuti stars—as a novel, independent pathway to discover dormant black holes in wide binaries, thereby opening a new observational channel in the quest to unveil the hidden compact-object population.
Key questions regarding the exact black hole mass, the detailed evolutionary pathway leading to such an extreme configuration, and the system’s ultimate fate remain. Future multi-wavelength observations, astrometric monitoring, and theoretical modeling of supernova kicks in binary systems will be essential to fully understand the origins and implications of this remarkable cosmic benchmark.
Materials & Correspondence
Correspondence and requests for materials should be addressed to Jia-Shu Niu. (email: jsniu@sxu.edu.cn).
H.F.X. acknowledges support from the National Natural Science Foundation of China (NSFC; No. 12303036) and the Scientific and Technological Innovation Programs of Higher Education Institutions in Shanxi (STIP; 2025Q032). All the authors acknowledge the TESS Science team and everyone who has contributed to making the TESS mission possible. We also acknowledge with thanks the variable star observations from the AAVSO International Database contributed by observers worldwide and used in this research. All the TESS data used in this paper can be found in MAST: http://dx.doi.org/10.17909/t9-nmc8-f686 (catalog 10.17909/t9-nmc8-f686).
{contribution}
J.S.N. and H.F.X. designed the research plan, performed the calculation, and wrote the manuscript. Y.Z. reviewed relevant literature, collected the TML, performed the frequency analysis, and wrote the initial version of Sec. 1 and part of Sec. 2.
Competing Interests
The authors declare that they have no competing financial interests.
Appendix A Methods
A.1 Photometric Data and Frequency Analysis
BE Lyn was observed by the Transiting Exoplanet Survey Satellite (TESS) in Sector 21 (January 2020) with a 120-second cadence. We retrieved the Presearch Data Conditioning Simple Aperture Photometry (PDCSAP) flux from the Mikulski Archive for Space Telescopes (MAST) and converted it to magnitudes using the TESS magnitude system (J. C. Smith et al., 2012; M. C. Stumpe et al., 2012, 2014; K. G. Stassun et al., 2019).
We extracted pulsation frequencies using an iterative prewhitening procedure (J.-S. Niu & H.-F. Xue, 2022; J.-S. Niu et al., 2023; J.-S. Niu & H.-F. Xue, 2024). The light curve was modeled as a superposition of sinusoids:
$m(t)=m_{0}+\sum_{i}a_{i}\sin[2\pi(f_{i}t+\phi_{i})],$ (A1)
where $m_{0}$ is the mean magnitude, and $a_{i}$, $f_{i}$, and $\phi_{i}$ are the amplitude, frequency, and phase of each component. Frequencies were accepted as significant if their signal-to-noise ratio (S/N) exceeded 5.6 (J.-S. Niu & H.-F. Xue, 2022; J.-S. Niu et al., 2023; J.-S. Niu & H.-F. Xue, 2024). After removing aliases, we identified 37 significant peaks, consisting of 7 independent modes and 30 harmonics or linear combinations (Table A1).
A.2 Compilation and Processing of Times of Maximum Light
The light curve is dominated by the fundamental mode ($f_{0}$), with the first overtone ($f_{1}$) contributing less than 3% of its amplitude (W. Xue et al., 2023). This allows the classical OCO-C method (C. Sterken, 2005) to be applied without significant contamination from multi-periodic effects.
We compiled times of maximum light (TML) from three sources: (1) historical literature, (2) the American Association of Variable Star Observers (AAVSO) database, and (3) TESS Sector 21 photometry. For AAVSO and TESS data, we determined TML by fitting a cubic polynomial to each light-curve maximum, with uncertainties estimated via Monte Carlo simulations.
To ensure temporal consistency, all times were converted to Barycentric Julian Date in the TDB system (BJDTDB). Heliocentric Julian Dates (HJD) were transformed using the tool of J. Eastman et al. (2010), while Julian Dates (JD) were converted using the Astropy time package.
Our final dataset includes 445 TML spanning 39 years (1986–2025), after excluding 13 outliers identified in a preliminary OCO-C analysis. For TML from C. Boonyarak et al. (2011) published without uncertainties, we adopted a conservative error of 0.0001 d, consistent with their reported precision and twice the mean uncertainty of our TESS-based measurements.
A.3 Orbital Modeling via MCMC
The OCO-C residuals were modeled with a function combining a quadratic term (intrinsic period variation) and a light-travel time effect (LTTE) term for orbital motion (H.-F. Xue & J.-S. Niu, 2020):
$C=\mathrm{BJD}_{0}+P_{0}E+\frac{1}{2}\beta E^{2}+A\left[\sqrt{1-e^{2}}\sin{\phi}\cos{\omega}+\cos{\phi}\sin{\omega}\right],$ (A2)
where $\phi$ is the eccentric anomaly from Kepler’s equation, $\mathrm{BJD}_{0}$ is the reference epoch, $P_{0}$ the pulsation period, $\beta$ the linear period variation rate, $A\equiv a\sin i/c$ the projected semi-major axis, $e$ the orbital eccentricity, $\omega$ the argument of periastron, $P_{\mathrm{orb}}$ the orbital period, and $T_{0}$ the time of periastron passage.
We employed a Markov Chain Monte Carlo (MCMC) approach using the emcee package (D. Foreman-Mackey et al., 2013) to sample the posterior distributions of all parameters. After chain convergence, we adopted the median of each posterior as the best-fit value, with uncertainties derived from the 5th and 95th percentiles. The best-fit model ($\chi^{2}/\mathrm{d.o.f.}=35.23$) is shown in Figure 1, with parameters listed in Table 1.
A.4 Asteroseismic Modeling of the Primary Star
We estimated the primary star’s properties by matching the observed pulsation characteristics to stellar models. Using MESA (Modules for Experiments in Stellar Astrophysics) (B. Paxton et al., 2011, 2013, 2015, 2018, 2019; A. S. Jermyn et al., 2023), we computed a grid of evolutionary tracks from the pre-main sequence to the red giant branch, with masses from 1.3 to 2.3 $M_{\odot}$ in steps of 0.01 $M_{\odot}$ and metallicity $Z=0.0034$ (from APOGEE-2 (Abdurro’uf et al., 2022), corresponding to [Fe/H] = -0.598). At each time step, we calculated adiabatic oscillation frequencies using GYRE (R. H. D. Townsend & S. A. Teitler, 2013; R. H. D. Townsend et al., 2018; J. Goldstein & R. H. D. Townsend, 2020; M. Sun et al., 2023). More details about the theoretical calculation can be referred to in J.-S. Niu & H.-F. Xue (2022) and J.-S. Niu & H.-F. Xue (2025).
The optimal model was selected by minimizing $\chi^{2}$ across three observables: the fundamental frequency ($f_{0}$), first-overtone frequency ($f_{1}$), and the intrinsic period variation rate $(1/P_{0})(dP_{0}/dt)$. This methodology follows established asteroseismic analyses of HADS (J.-S. Niu et al., 2017; H.-F. Xue et al., 2018, 2022). The best-fit model parameters are given in Table A2.
A.5 Dynamical Constraints on the Companion
The companion’s nature was inferred from the mass function,
$f(M_{2})=\frac{(M_{2}\sin i)^{3}}{(M_{1}+M_{2})^{2}}=\frac{4\pi^{2}}{GP_{\mathrm{orb}}^{2}}(a\sin i)^{3},$ (A3)
combined with the constraints that the periastron separation $d_{\mathrm{peri}}=a(1-e)(M_{1}+M_{2})/M_{2}$ must exceed the primary’s radius $R_{1}$, and the Roche lobe radius $R_{1,\mathrm{R}}$ must exceed the primary’s radius $R_{1}$. The Roche lobe radius was calculated using the approximation of P. P. Eggleton (1983). This yields an upper limit on the inclination ($i\lesssim 4.0^{\circ}$) and a corresponding lower limit on the companion mass ($M_{2}\gtrsim 17.5M_{\odot}$), as shown in Figure 3.
Appendix B Additional Figures
Figure A1: Light curve of BE Lyn from TESS Sector 21. The lower panel shows a one-day zoom for clarity.
Figure A2: Frequency spectrum of BE Lyn. The main panel (0100c/d100\ $\mathrm${c/d}) marks the fundamental ($f_{0}$, red) and first-overtone ($f_{1}$, green) modes. The inset highlights the low-amplitude, non-radial mode region ($\sim 9-17\ \mathrm{c/d}$) with a linear y-axis.
Appendix C Additional Tables
Table A1: Frequency solution for BE Lyn from TESS Sector 21.
NO. f(c/d)f\ ($\mathrm${c/d}) $\sigma_{f}\ (\mathrm{c/d})$ a(mmag)a\ ($\mathrm${mmag}) $\sigma_{a}\ (\mathrm{mmag})$ S/N Marks
F1 10.4309 0.0005 116 3 42.4 $f_{0}$
F2 20.8618 0.0005 38.9 0.9 42.1 $2f_{0}$
F3 31.2926 0.0005 14.3 0.3 42.4 $3f_{0}$
F4 41.7236 0.0005 6.0 0.1 44.2 $4f_{0}$
F5 13.4226 0.0005 2.59 0.07 37.8 $f_{1}$
F6 52.1544 0.0005 2.57 0.06 40.3 $5f_{0}$
F7 62.5852 0.0005 1.41 0.03 41.3 $6f_{0}$
F8 23.8534 0.0005 1.04 0.03 38.4 $f_{0}+f_{1}$
F9 73.0156 0.0006 0.79 0.02 34.6 $7f_{0}$
F10 2.9914 0.0007 0.71 0.03 27.4 $f_{0}+f_{1}$
F11 34.2846 0.0008 0.50 0.02 27.0 $2f_{0}+f_{1}$
F13 83.4470 0.0006 0.47 0.01 31.8 $8f_{0}$
F14 7.440 0.001 0.37 0.02 21.3 $2f_{0}-f_{1}$
F15 44.7150 0.0008 0.29 0.01 24.4 $3f_{0}+f_{1}$
F16 93.8765 0.0009 0.26 0.01 23.8 $9f_{0}$
F19 55.147 0.001 0.16 0.01 17.7 $4f_{0}+f_{1}$
F20 9.701 0.003 0.15 0.02 7.2 $f_{2}$
F22 104.309 0.002 0.14 0.01 14.0 $10f_{0}$
F24 17.869 0.002 0.11 0.01 10.5 $3f_{0}-f_{1}$
F25 65.579 0.002 0.10 0.01 10.4 $5f_{0}+f_{1}$
F26 10.684 0.003 0.10 0.02 6.0 $f_{3}$
F27 9.377 0.003 0.09 0.01 6.0 $f_{4}$
F28 114.740 0.002 0.076 0.008 9.8 $11f_{0}$
F29 14.009 0.003 0.08 0.01 6.4 $f_{5}$
F30 19.810 0.003 0.08 0.01 6.2 $f_{0}+f_{4}$
F31 20.133 0.003 0.07 0.01 5.9 $f_{0}+f_{2}$
F32 21.112 0.004 0.070 0.01 5.7 $f_{0}+2f_{3}$
F33 76.006 0.003 0.067 0.009 7.5 $6f_{0}+f_{1}$
F34 16.336 0.003 0.07 0.01 6.8 $f_{6}$
F35 28.299 0.003 0.058 0.009 6.1 $4f_{0}-f_{1}$
F36 38.729 0.003 0.052 0.008 6.1 $5f_{0}-f_{1}$
F37 86.438 0.004 0.045 0.008 5.8 $7f_{0}+f_{1}$
Table A2: Parameters of the best-fit seismic model for BE Lyn’s primary star.
Parameter Best-fit Model Observed
$f_{0}$ (c/d) 10.4288 $10.4309\pm 0.0005$
$f_{1}$ (c/d) 13.4243 $13.4226\pm 0.0005$
$(1/P_{0})(dP_{0}/dt)$ (yr-1) $1.34\times 10^{-8}$ $(1.02^{+0.08}_{-0.08})\times 10^{-8}$
$\log T_{\mathrm{eff}}$ 3.902
$\log(L/L_{\odot})$ 1.326
$\log g$ 3.865
$\log(R/R_{\odot})$ 0.382
Age (Gyr) 1.59
Mass ($M_{\odot}$) 1.55
[…]
Table A3: Times of maximum light of BE Lyn. The numbers in the fourth, eighth, and twelfth columns indicate the different data sources: (1) T. Oja (1987); (2) E. Rodríguez et al. (1990); (3) Y. Liu et al. (1991); (4) E. Wunder et al. (1992); (5) Q. Tang et al. (1992); (6) Z. Liu & S. Jiang (1994); (7) L. L. Kiss & K. Szatmary (1995); (8) A. Derekas et al. (2003); (9) J. N. Fu & S. Y. Jiang (2005); (10) R. Szakats et al. (2008); (11) J. Hubscher (2005);(12) J. Hubscher et al. (2005); (13) J. Hubscher et al. (2006); (14) J. Hübscher et al. (2005); (15) G. Klingenberg et al. (2006); (16) C. Boonyarak et al. (2011); (17) J. Hubscher (2007); (18) J. Hubscher et al. (2009); (19) AAVSO; (20) J. Hubscher et al. (2013); (21) J. H. Peña et al. (2015); (22) J. Hubscher (2014); (23) J. Hübscher (2017); (24) TESS; (25) L. Pagel (2022); (26) L. Pagel (2021); (27) L. Pagel (2023).
TML($\mathrm{BJD}$) Error Detector Source TML($\mathrm{BJD}$) Error Detector Source TML($\mathrm{BJD}$) Error Detector Source
2446498.3387 0.0002 pe 1 2454433.6515 0.0008 CCD 18 2458882.47489 0.00004 CCD 24
2446507.3500 0.0002 pe 1 2454837.356aaThese TML are observed without filter. 0.001 CCD 18 2458882.57111 0.00004 CCD 24
2446507.4466 0.0002 pe 1 2454858.6447**These TML are excluded in the global fitting because of the significant deviations in the preliminary OCO-C analysis. 0.0001 CCD 16 2458882.66680 0.00003 CCD 24
2446508.4052 0.0002 pe 1 2454871.3891 0.0001 CCD 16 2458882.76233 0.00004 CCD 24
2446509.4600 0.0002 pe 1 2454946.3599 0.0001 CCD 16 2458882.85851 0.00004 CCD 24
2446510.4184 0.0002 pe 1 2455309.6105 0.0002 CCD 19 2458882.95462 0.00004 CCD 24
2446524.4151 0.0002 pe 1 2455646.3031 0.0006 CCD 20 2458883.04995 0.00004 CCD 24
2446950.4592 0.0001 pe 2 2455646.3992 0.0005 CCD 20 2458883.14582 0.00004 CCD 24
2446951.4185 0.0001 pe 2 2455968.3297 0.0007 CCD 20 2458883.24205 0.00004 CCD 24
2447115.6426 0.0001 pe 2 2456002.363 0.002 CCD 20 2458883.33778 0.00003 CCD 24
2447118.6144 0.0001 pe 2 2456046.465 0.002 CCD 20 2458883.43331 0.00003 CCD 24
2447118.7104 0.0002 pe 2 2456050.393 0.004 CCD 20 2458883.52961 0.00004 CCD 24
2447121.6822 0.0001 pe 2 2456347.780 0.002 CCD 21 2458883.62569 0.00003 CCD 24
2447219.5648 0.0002 pe 2 2456347.876aaThese TML are observed without filter. 0.002 CCD 21 2458883.72091 0.00003 CCD 24
2447265.1032 0.0001 pe 3 2456360.820 0.002 CCD 21 2458883.81699 0.00004 CCD 24
2447486.3699 0.0001 pe 3 2456371.366 0.003 DSLR 22 2458883.91322 0.00004 CCD 24
2447488.2859 0.0001 pe 3 2456374.815 0.002 CCD 21 2458886.98108 0.00003 CCD 24
2447488.3818 0.0001 pe 3 2456376.638 0.002 CCD 21 2458887.07648 0.00003 CCD 24
2447489.2457 0.0001 pe 3 2456376.733 0.002 CCD 21 2458887.17252 0.00004 CCD 24
2447489.3419 0.0001 pe 3 2456657.440bbThese TML are observed by IR cut-off filter. 0.001 CCD 23 2458887.26876 0.00004 CCD 24
2447493.1762 0.0001 pe 3 2457454.955**These TML are excluded in the global fitting because of the significant deviations in the preliminary OCO-C analysis. 0.002 DSLR 19 2458887.36421 0.00003 CCD 24
2447493.2710 0.0001 pe 3 2457458.3346 0.0007 CCD 23 2458887.45982 0.00003 CCD 24
2447542.2612 0.0001 pe 3 2457458.4306 0.0006 CCD 23 2458887.55606 0.00003 CCD 24
2447544.2737 0.0001 pe 3 2458847.3857ccThese TML are observed by the TG band (DSLR green-channel). 0.0005 DSLR 19 2458887.65191 0.00003 CCD 24
2447544.3702 0.0001 pe 3 2458847.4825ccThese TML are observed by the TG band (DSLR green-channel). 0.0003 DSLR 19 2458887.74747 0.00003 CCD 24
2447551.4654 0.0001 pe 2 2458847.5786ccThese TML are observed by the TG band (DSLR green-channel). 0.0003 DSLR 19 2458887.84355 0.00004 CCD 24
2447551.5613 0.0001 pe 2 2458847.6726ccThese TML are observed by the TG band (DSLR green-channel). 0.0005 DSLR 19 2458887.93986 0.00004 CCD 24
2447551.6567 0.0001 pe 2 2458847.7696ccThese TML are observed by the TG band (DSLR green-channel). 0.0002 DSLR 19 2458888.03535 0.00004 CCD 24
2447553.5743 0.0001 pe 2 2458870.49132 0.00004 CCD 24 2458888.13094 0.00004 CCD 24
2447553.6702 0.0001 pe 2 2458870.58827 0.00003 CCD 24 2458888.22723 0.00004 CCD 24
2447627.0095 0.0001 pe 3 2458870.68287 0.00004 CCD 24 2458888.32305 0.00003 CCD 24
2447634.0087 0.0001 pe 3 2458870.77880 0.00004 CCD 24 2458888.41851 0.00004 CCD 24
2448347.3735 0.0001 pe 4 2458870.87510 0.00003 CCD 24 2458888.51457 0.00004 CCD 24
2448347.4686 0.0001 pe 4 2458870.97065 0.00004 CCD 24 2458888.61086 0.00004 CCD 24
2448357.3451 0.0001 pe 4 2458871.06612 0.00004 CCD 24 2458888.70635 0.00003 CCD 24
2448358.3995 0.0001 pe 4 2458871.16239 0.00004 CCD 24 2458888.80202 0.00003 CCD 24
2448359.3564 0.0001 pe 4 2458871.25848 0.00004 CCD 24 2458888.89837 0.00004 CCD 24
2448367.5067 0.0001 pe 4 2458871.35387 0.00003 CCD 24 2458888.99415 0.00004 CCD 24
2448624.3407 0.0001 pe 4 2458871.44980 0.00004 CCD 24 2458889.08950 0.00004 CCD 24
2448707.0770 0.0001 pe 5 2458871.54610 0.00004 CCD 24 2458889.18563 0.00004 CCD 24
2448710.0480 0.0001 pe 5 2458871.64164 0.00004 CCD 24 2458889.28186 0.00004 CCD 24
2448710.1441 0.0001 pe 5 2458871.73743 0.00004 CCD 24 2458889.37750 0.00004 CCD 24
2448710.2397 0.0001 pe 5 2458871.83361 0.00004 CCD 24 2458889.47307 0.00004 CCD 24
2448715.1294 0.0001 pe 5 2458871.92941 0.00003 CCD 24 2458889.56947 0.00004 CCD 24
2448715.2260 0.0001 pe 5 2458872.02491 0.00003 CCD 24 2458889.66535 0.00004 CCD 24
2449018.0775 0.0001 pe 6 2458872.12093 0.00004 CCD 24 2458889.76080 0.00004 CCD 24
2449018.1733 0.0001 pe 6 2458872.21706 0.00004 CCD 24 2458889.85684 0.00003 CCD 24
2449018.2692 0.0001 pe 6 2458872.31280 0.00004 CCD 24 2458889.95305 0.00004 CCD 24
2449018.3651 0.0001 pe 6 2458872.40836 0.00003 CCD 24 2458890.04850 0.00003 CCD 24
2449019.0359 0.0001 pe 6 2458872.50464 0.00004 CCD 24 2458890.14419 0.00004 CCD 24
2449021.0507 0.0001 pe 6 2458872.60063 0.00003 CCD 24 2458890.24059 0.00004 CCD 24
2449022.0086 0.0001 pe 6 2458872.69606 0.00004 CCD 24 2458890.33633 0.00004 CCD 24
2449030.0612 0.0001 pe 6 2458872.79194 0.00004 CCD 24 2458890.43185 0.00004 CCD 24
2449031.0199 0.0001 pe 6 2458872.88832 0.00004 CCD 24 2458890.52798 0.00003 CCD 24
2449400.0222 0.0001 pe 6 2458872.98382 0.00003 CCD 24 2458890.62419 0.00004 CCD 24
2449400.1180 0.0001 pe 6 2458873.07934 0.00004 CCD 24 2458890.71956 0.00003 CCD 24
2449401.0760 0.0001 pe 6 2458873.17571 0.00004 CCD 24 2458890.81532 0.00003 CCD 24
2449401.1720 0.0001 pe 6 2458873.27175 0.00004 CCD 24 2458890.91166 0.00004 CCD 24
2449413.0600 0.0001 pe 6 2458873.36714 0.00004 CCD 24 2458891.00736 0.00003 CCD 24
2449413.1560 0.0001 pe 6 2458873.46306 0.00004 CCD 24 2458891.10285 0.00003 CCD 24
2449416.1279 0.0001 pe 6 2458873.55933 0.00004 CCD 24 2458891.19901 0.00004 CCD 24
2449749.4657 0.0002 pe 7 2458873.65484 0.00003 CCD 24 2458891.29526 0.00004 CCD 24
2449749.5616 0.0002 pe 7 2458873.75062 0.00003 CCD 24 2458891.39060 0.00003 CCD 24
2449749.6575 0.0002 pe 7 2458873.84678 0.00004 CCD 24 2458891.48640 0.00004 CCD 24
2449754.3549 0.0004 pe 7 2458873.94276 0.00003 CCD 24 2458891.58276 0.00004 CCD 24
2449762.4077 0.0004 pe 8 2458874.03815 0.00004 CCD 24 2458891.67854 0.00004 CCD 24
2450114.5373 0.0005 pe 8 2458874.13411 0.00004 CCD 24 2458891.77394 0.00004 CCD 24
2450139.4632 0.0004 pe 8 2458874.23035 0.00004 CCD 24 2458891.87012 0.00004 CCD 24
2450140.4216 0.0002 pe 8 2458874.32609 0.00003 CCD 24 2458891.96628 0.00004 CCD 24
2450500.5078 0.0002 pe 8 2458874.42166 0.00003 CCD 24 2458892.06157 0.00003 CCD 24
2450844.3916 0.0001 pe 8 2458874.51795 0.00004 CCD 24 2458892.15742 0.00003 CCD 24
2450844.4874 0.0001 pe 8 2458874.61393 0.00004 CCD 24 2458892.25382 0.00004 CCD 24
2451205.3404 0.0002 pe 9 2458874.70925 0.00004 CCD 24 2458892.34951 0.00004 CCD 24
2451205.5322 0.0001 pe 8 2458874.80541 0.00004 CCD 24 2458892.44507 0.00004 CCD 24
2451205.6280 0.0001 pe 8 2458874.90155 0.00004 CCD 24 2458892.54120 0.00003 CCD 24
2451205.7239 0.0001 pe 8 2458874.99696 0.00004 CCD 24 2458892.63732 0.00003 CCD 24
2451207.3549 0.0001 pe 9 2458875.09267 0.00003 CCD 24 2458892.73273 0.00004 CCD 24
2451209.2717 0.0001 pe 9 2458875.18902 0.00004 CCD 24 2458892.82861 0.00004 CCD 24
2451209.3669 0.0001 pe 9 2458875.28492 0.00003 CCD 24 2458892.92495 0.00004 CCD 24
2451237.2645 0.0002 pe 8 2458875.38040 0.00004 CCD 24 2458893.02060 0.00004 CCD 24
2451237.3603 0.0002 pe 8 2458875.47644 0.00003 CCD 24 2458893.11598 0.00004 CCD 24
2451237.4562 0.0002 pe 8 2458875.57261 0.00003 CCD 24 2458893.21221 0.00005 CCD 24
2451237.5521 0.0002 pe 8 2458875.66818 0.00003 CCD 24 2458893.30843 0.00004 CCD 24
2451239.3737 0.0003 pe 8 2458875.76375 0.00004 CCD 24 2458893.40388 0.00003 CCD 24
2451239.4696 0.0003 pe 8 2458875.85999 0.00004 CCD 24 2458893.49967 0.00004 CCD 24
2451240.4287 0.0001 pe 8 2458875.95596 0.00004 CCD 24 2458893.59602 0.00004 CCD 24
2451240.5246 0.0001 pe 8 2458876.05147 0.00003 CCD 24 2458893.69190 0.00004 CCD 24
2451240.6205 0.0001 pe 8 2458876.14746 0.00004 CCD 24 2458893.78726 0.00004 CCD 24
2451243.3041 0.0002 pe 8 2458876.24366 0.00004 CCD 24 2458893.88343 0.00004 CCD 24
2451243.4000 0.0002 pe 8 2458876.33924 0.00003 CCD 24 2458893.97944 0.00004 CCD 24
2451246.2764 0.0001 pe 8 2458876.43493 0.00004 CCD 24 2458894.07487 0.00004 CCD 24
2451246.3723 0.0001 pe 8 2458876.53122 0.00004 CCD 24 2458894.17079 0.00004 CCD 24
2451246.4681 0.0001 pe 8 2458876.62707 0.00003 CCD 24 2458894.26713 0.00004 CCD 24
2451293.3486 0.0001 pe 8 2458876.72311 0.00006 CCD 24 2458894.36280 0.00004 CCD 24
2451293.4444 0.0001 pe 8 2458876.81860 0.00004 CCD 24 2458894.45822 0.00004 CCD 24
2451311.4677 0.0002 pe 8 2458876.91469 0.00004 CCD 24 2458894.55453 0.00004 CCD 24
2451313.4810 0.0002 pe 8 2458877.01031 0.00003 CCD 24 2458894.65061 0.00004 CCD 24
2451585.3671 0.0002 pe 8 2458877.10601 0.00004 CCD 24 2458894.74601 0.00003 CCD 24
2451929.4429 0.0002 pe 8 2458877.20231 0.00004 CCD 24 2458894.84178 0.00003 CCD 24
2451929.5387 0.0002 pe 8 2458877.29814 0.00004 CCD 24 2458894.93818 0.00004 CCD 24
2451965.2983 0.0001 CCD 8 2458877.39364 0.00003 CCD 24 2458895.03386 0.00004 CCD 24
2451965.3941 0.0001 CCD 8 2458877.48958 0.00004 CCD 24 2458895.12929 0.00004 CCD 24
2451965.4903 0.0001 CCD 8 2458877.58585 0.00003 CCD 24 2458895.22565 0.00004 CCD 24
2451985.3347 0.0002 CCD 8 2458877.68135 0.00003 CCD 24 2458895.32171 0.00003 CCD 24
2451985.5264 0.0002 CCD 8 2458877.77708 0.00004 CCD 24 2458895.41708 0.00004 CCD 24
2452252.5198**These TML are excluded in the global fitting because of the significant deviations in the preliminary OCO-C analysis. 0.0001 CCD 8 2458877.87348 0.00004 CCD 24 2458895.51297 0.00004 CCD 24
2452371.4010 0.0002 CCD 8 2458877.96917 0.00004 CCD 24 2458895.60936 0.00004 CCD 24
2452371.4969 0.0002 CCD 8 2458878.06464 0.00004 CCD 24 2458895.70493 0.00004 CCD 24
2452396.3273 0.0001 CCD 8 2458878.16078 0.00004 CCD 24 2458895.80042 0.00004 CCD 24
2452396.4232 0.0001 CCD 8 2458878.25694 0.00004 CCD 24 2458895.89673 0.00004 CCD 24
2452981.5147 0.0001 CCD 10 2458878.35252 0.00004 CCD 24 2458895.99269 0.00004 CCD 24
2452981.6106 0.0001 CCD 10 2458878.44826 0.00004 CCD 24 2458896.08806 0.00004 CCD 24
2453056.485aaThese TML are observed without filter. 0.001 pe 11 2458878.54457 0.00004 CCD 24 2458896.18406 0.00004 CCD 24
2453056.582aaThese TML are observed without filter. 0.001 pe 11 2458878.64034 0.00003 CCD 24 2458896.28041 0.00004 CCD 24
2453094.3522 0.0008 pe 12 2458878.73583 0.00003 CCD 24 2458896.37596 0.00004 CCD 24
2453094.371b*b*footnotemark: 0.007 pe 11 2458878.83192 0.00003 CCD 24 2458896.47141 0.00004 CCD 24
2453094.4474 0.0008 pe 12 2458878.92801 0.00004 CCD 24 2458896.56775 0.00004 CCD 24
2453111.323bbThese TML are observed by IR cut-off filter. 0.002 pe 11 2458879.02343 0.00004 CCD 24 2458896.66375 0.00004 CCD 24
2453111.419bbThese TML are observed by IR cut-off filter. 0.003 pe 11 2458879.11923 0.00004 CCD 24 2458896.75910 0.00004 CCD 24
2453111.517b*b*footnotemark: 0.003 pe 11 2458879.21571 0.00004 CCD 24 2458896.85515 0.00004 CCD 24
2453118.4162 0.0001 CCD 10 2458879.31139 0.00004 CCD 24 2458896.95144 0.00004 CCD 24
2453349.462 0.001 CCD 13 2458879.40675 0.00004 CCD 24 2458897.04707 0.00003 CCD 24
2453349.558 0.001 CCD 13 2458879.50304 0.00004 CCD 24 2458897.14258 0.00003 CCD 24
2453349.653 0.001 CCD 13 2458879.59920 0.00004 CCD 24 2458897.23891 0.00004 CCD 24
2453352.439**These TML are excluded in the global fitting because of the significant deviations in the preliminary OCO-C analysis. 0.001 vis 14 2458879.69459 0.00003 CCD 24 2458897.33481 0.00003 CCD 24
2453410.542b*b*footnotemark: 0.005 pe 12 2458879.79036 0.00003 CCD 24 2458897.43031 0.00004 CCD 24
2453416.2791a*a*footnotemark: 0.0005 CCD 15 2458879.88672 0.00004 CCD 24 2458897.52626 0.00004 CCD 24
2453416.3759aaThese TML are observed without filter. 0.0005 CCD 15 2458879.98240 0.00003 CCD 24 2458897.62254 0.00004 CCD 24
2453443.6034aaThese TML are observed without filter. 0.0004 CCD 15 2458880.07791 0.00003 CCD 24 2458897.71815 0.00004 CCD 24
2453472.372b*b*footnotemark: 0.005 pe 12 2458880.17417 0.00004 CCD 24 2458954.378ccThese TML are observed by the TG band (DSLR green-channel). 0.003 CCD 25
2453472.465b*b*footnotemark: 0.002 pe 12 2458880.27026 0.00004 CCD 24 2458955.35c*c*footnotemark: 0.01 DSLR 26
2453789.1190 0.0001 CCD 16 2458880.36559 0.00003 CCD 24 2458961.37666ccThese TML are observed by the TG band (DSLR green-channel). 0.00009 DSLR 19
2453789.2140 0.0001 CCD 16 2458880.46137 0.00003 CCD 24 2458961.462**These TML are excluded in the global fitting because of the significant deviations in the preliminary OCO-C analysis. 0.001 DSLR 26
2453790.0752 0.0001 CCD 16 2458880.55773 0.00004 CCD 24 2458965.403ccThese TML are observed by the TG band (DSLR green-channel). 0.001 CCD 25
2453791.0344 0.0001 CCD 16 2458880.65346 0.00004 CCD 24 2458966.5566c*c*footnotemark: 0.0001 DSLR 19
2453791.1307 0.0001 CCD 16 2458880.74899 0.00003 CCD 24 2458967.51176ccThese TML are observed by the TG band (DSLR green-channel). 0.00009 DSLR 19
2453792.0897 0.0001 CCD 16 2458880.84519 0.00004 CCD 24 2458975.4703ccThese TML are observed by the TG band (DSLR green-channel). 0.0003 DSLR 19
2453792.1859 0.0001 CCD 16 2458880.94130 0.00003 CCD 24 2459277.5562**These TML are excluded in the global fitting because of the significant deviations in the preliminary OCO-C analysis. 0.0002 CCD 19
2453793.0473 0.0001 CCD 16 2458881.03670 0.00004 CCD 24 2459297.399ccThese TML are observed by the TG band (DSLR green-channel). 0.004 DSLR 25
2453793.1448 0.0001 CCD 16 2458881.13260 0.00004 CCD 24 2459303.345c*c*footnotemark: 0.004 DSLR 25
2453798.3189a*a*footnotemark: 0.0002 CCD 15 2458881.22889 0.00004 CCD 24 2459689.410ccThese TML are observed by the TG band (DSLR green-channel). 0.004 DSLR 27
2454027.5469 0.0001 CCD 10 2458881.32458 0.00004 CCD 24 2459690.3667ccThese TML are observed by the TG band (DSLR green-channel). 0.0004 DSLR 19
2454027.6428 0.0001 CCD 10 2458881.42006 0.00003 CCD 24 2459990.4390ccThese TML are observed by the TG band (DSLR green-channel). 0.0008 CCD 19
2454028.5056 0.0001 CCD 10 2458881.51638 0.00004 CCD 24 2459990.5347ccThese TML are observed by the TG band (DSLR green-channel). 0.0009 CCD 19
2454028.6015 0.0001 CCD 10 2458881.61236 0.00003 CCD 24 2460049.3993ccThese TML are observed by the TG band (DSLR green-channel). 0.0001 DSLR 19
2454036.4631 0.0002 CCD 10 2458881.70772 0.00004 CCD 24 2460254.8461ccThese TML are observed by the TG band (DSLR green-channel). 0.0002 DSLR 19
2454036.5588 0.0002 CCD 10 2458881.80368 0.00004 CCD 24 2460700.640ccThese TML are observed by the TG band (DSLR green-channel). 0.001 CCD 19
2454036.6546 0.0002 CCD 10 2458881.89996 0.00004 CCD 24 2460708.5016ccThese TML are observed by the TG band (DSLR green-channel). 0.0007 CCD 19
2454039.4346 0.0001 CCD 10 2458881.99566 0.00003 CCD 24 2460714.542ccThese TML are observed by the TG band (DSLR green-channel). 0.001 CCD 19
2454039.5305 0.0001 CCD 10 2458882.09110 0.00004 CCD 24 2460783.5680ccThese TML are observed by the TG band (DSLR green-channel). 0.0006 DSLR 19
2454039.6263 0.0001 CCD 10 2458882.18743 0.00004 CCD 24 2460783.6660ccThese TML are observed by the TG band (DSLR green-channel). 0.0007 DSLR 19
2454221.3950 0.0009 CCD 17 2458882.28342 0.00003 CCD 24 2460805.4251ccThese TML are observed by the TG band (DSLR green-channel). 0.0002 DSLR 19
2454222.4498 0.0008 CCD 17 2458882.37876 0.00004 CCD 24
Note. —

### v2

@@
-BE Lyncis: An Extremely Eccentric Binary with the Nearest Known Black Hole
+BE Lyncis: A Pulsating Star in the Most Eccentric Binary with a Massive Unseen Companion
 Abstract
-We report the discovery of an exceptionally eccentric binary system, BE Lyncis (BE Lyn), which hosts the nearest known black hole (BH) to Earth.
-Through the analysis of TESS photometry combined with an extensive set of times of maximum light spanning 39 years, we identify BE Lyn as a high-amplitude $\delta$ Scuti star in a binary with an orbital period of $\approx 15.9$ years and an extraordinary orbital eccentricity of $e=0.9989^{+0.0008}_{-0.0021}$ (>0.9968>0.9968 at 95% confidence)—the highest reliably measured for any binary system.
-Dynamical constraints impose an upper limit on the orbital inclination of $i\lesssim 4.0^{\circ}$, corresponding to a companion mass of $M_{2}\gtrsim 17.5~M_{\odot}$, which unequivocally favors a black hole.
-This system provides a unique laboratory for studying asteroseismology in strong gravitational fields, the formation of black holes via asymmetric supernovae, and the evolution of extreme binary systems.
-Our work demonstrates, for the first time, the successful application of the light-travel time effect in a pulsating variable to unveil a dormant black hole, establishing a novel method for BH detection in non-interacting binaries.
+We report the discovery of an exceptionally eccentric binary system, BE Lyncis (BE Lyn), which hosts a compact companion with mass $\gtrsim 2.5~M_{\odot}$.
+By combining TESS photometry with an extensive set of times of maximum light spanning 39 years, we identify BE Lyn as a high-amplitude $\delta$ Scuti star in a binary with an orbital period of $\approx 15.9$ years and an extraordinary eccentricity of $e=0.9989^{+0.0008}_{-0.0021}$ (>0.9968>0.9968 at 95% confidence)—the most extreme eccentricity reliably measured for any binary system.
+Dynamical constraints limit the orbital inclination to $i\lesssim 10.1^{\circ}$, implying a companion mass $M_{2}\gtrsim 2.5~M_{\odot}$, which spans the mass gap between neutron stars and black holes.
+If the companion is a neutron star, it would be the most massive known; if a black hole, it would be the closest to Earth.
+This system provides a unique laboratory for studying asteroseismology in strong gravitational fields, as well as the formation and evolution of extremely eccentric binaries.
+Our work demonstrates the use of the light-travel time effect in a pulsating star to reveal a compact companion, offering a novel method for detecting black holes in non-interacting binaries.
+software: astropy (Astropy Collaboration et al., 2022), Lightkurve (Lightkurve Collaboration et al., 2018), NumPy (C.
+R.
+Harris et al., 2020), SciPy (P.
+Virtanen et al., 2020), matplotlib (J.
+D.
+Hunter, 2007), emcee (D.
+Foreman-Mackey et al., 2013), MESA (B.
+Paxton et al., 2011, 2013, 2015, 2018, 2019; A.
+S.
+Jermyn et al., 2023), GYRE (R.
+H.
+D.
+Townsend & S.
+A.
+Teitler, 2013)
 I Introduction
-Black holes (BHs) of stellar origin represent one of the most enigmatic predictions of general relativity.
-While several dozen stellar-mass BHs have been identified in X-ray binaries through accretion signatures (R.
-W.
-Romani, 1998), the population of dormant (non-accreting) BHs in detached binaries remains largely unexplored.
-Detecting such systems is challenging, as they typically reveal themselves only through subtle gravitational interactions with luminous companions.
-Recent astrometric surveys, such as Gaia, have uncovered a handful of such quiescent BH binaries (K.
-El-Badry et al., 2023; A.
-Tanikawa et al., 2023; Gaia Collaboration et al., 2024), yet they remain exceedingly rare.
-An independent, complementary detection method—leveraging the precise clocks provided by pulsating stars—offers a promising pathway to expand this census and probe different regions of parameter space.
-Pulsating $\delta$ Scuti stars are particularly well-suited for this purpose.
-These stars, with spectral types A–F, occupy the intersection of the main sequence and the classical Cepheid instability strip in the Hertzsprung–Russell diagram.
-Their oscillations, driven by the $\kappa$ mechanism operating in the partial ionization zone of helium, provide remarkably stable periodic signals with periods ranging from 15 minutes to 8 hours (T.
+Pulsating $\delta$ Scuti stars, with spectral types A–F, occupy the intersection of the main sequence and the classical Cepheid instability strip in the Hertzsprung–Russell diagram.
+Their oscillations, driven by the $\kappa$ mechanism operating in the helium partial ionization zone, provide stable periodic signals with periods ranging from 15 minutes to 8 hours (T.
 Kallinger et al., 2008; G.
@@
 Steindl et al., 2022).
-The subclass of high-amplitude $\delta$ Scuti stars (HADS) exhibits photometric variations exceeding 0.1 mag and typically rotates slowly ($v\sin i\leq 30$ km s-1), making them ideal precision tools for detecting the minite period modulations induced by orbital motion in binary systems—the light-travel time effect (LTTE).
-BE Lyncis (BE Lyn; HD 79889) is a well-studied HADS with a mean magnitude of $\langle V\rangle=8.8$ mag, spectral type A3, and sub-solar metallicity ([Fe/H] = -0.598) (C.
+High-amplitude $\delta$ Scuti stars (HADS), a subclass exhibiting photometric variations exceeding 0.1 mag and typically slow rotation ($v\sin i\leq 30$ km s-1), serve as precision celestial clocks.
+Their stability makes them ideally suited for detecting the minute period modulations induced by orbital motion in binary systems—a phenomenon known as the light-travel time effect (LTTE).
+BE Lyncis (BE Lyn; HD 79889) is a well-studied HADS with a mean magnitude of $\langle V\rangle=8.8$ mag, spectral type A3, and sub-solar metallicity ([Fe/H] = 0.598-0.598) (C.
 Kim et al., 2012; Abdurro’uf et al., 2022).
-It was first identified as a variable star by T.
+First identified as a variable star by T.
 Oja (1986) and later confirmed as a HADS with a pulsation period of $\sim 0.09587$ days (T.
-Oja, 1987).
-Subsequent studies of its period variations revealed anomalies in OCO-C diagrams, occasionally suggestive of a parabolic trend indicative of a binary companion (e.g., E.
+Oja, 1987), it has been the subject of numerous period variation studies.
+Anomalies in its OCO-C diagram have long hinted at the presence of a binary companion, occasionally manifesting as a parabolic trend (e.g., E.
 Rodriguez et al., 1990; Y.
@@
 Liakos & P.
-Niarchos, 2017), though inferred orbital parameters varied significantly due to limitations in the temporal coverage and precision of the data.
+Niarchos, 2017), though the inferred orbital parameters varied substantially owing to limited temporal coverage and data precision.
 Two key developments now enable a definitive reassessment of this system.
-First, high-precision, continuous photometry from the Transiting Exoplanet Survey Satellite (TESS) allows for an exquisite characterization of the primary star’s pulsation modes and yields significantly more accurate times of maximum light.
+First, high-precision, continuous photometry from the Transiting Exoplanet Survey Satellite (TESS) allows exquisite characterization of the primary star’s pulsation modes and yields significantly more accurate times of maximum light.
 Second, modern computational techniques, such as robust multi-parameter global fitting via Markov Chain Monte Carlo (MCMC) methods, permit a physically consistent and precise determination of binary orbital parameters.
-Here, we present a comprehensive analysis of BE Lyn, combining TESS photometry with a newly compiled, 39-year baseline of times of maximum light.
-Our work not only confirms its binary nature but reveals an astrophysical gem: an ultra-wide, extremely eccentric orbit harboring a massive black hole companion.
-This system holds multiple distinctions: (i) it possesses the highest orbital eccentricity ever reliably measured in a binary; (ii) it hosts the closest known black hole to Earth, at a distance of approximately 250 pc (C.
+Here we present a comprehensive analysis of BE Lyn, combining TESS photometry with a newly compiled, 39-year baseline of times of maximum light.
+Our work not only confirms its binary nature but reveals an astrophysical gem: an ultra-wide, extremely eccentric orbit harbouring a compact companion of mass $\gtrsim 2.5~M_{\odot}$.
+This places the companion spanning the mass gap between neutron stars and black holes.
+If it is a neutron star, it would be the most massive known; if a black hole, it would be the closest to Earth, at a distance of approximately 250 pc (C.
 A.
 L.
-Bailer-Jones et al., 2021); (iii) it is the only known binary comprising a pulsating star and a quiescent black hole, opening a new window for asteroseismology in strong gravity; and (iv) its extreme orbital configuration places stringent constraints on the natal kick imparted during the black hole’s formation, challenging models of asymmetric supernova explosions.
-The detection methodology itself—using pulsation timing in a $\delta$ Scuti star—pioneers a new, independent channel for discovering dormant black holes.
+Bailer-Jones et al., 2021).
+The system holds multiple distinctions: (i) it possesses the highest orbital eccentricity ever reliably measured in any binary; (ii) it is the only known binary comprising a pulsating star and a massive compact companion, opening a new window for asteroseismology in strong gravitational fields; (iii) its extreme eccentricity ($e=0.9989$) provides a stringent observational test for models of supernova natal kicks and post-supernova mass transfer, offering critical insights into the formation pathways of such extreme binaries; and (iv) it validates the LTTE method using pulsating stars as a novel, independent pathway for discovering quiescent compact objects in non-interacting binaries.
+BE Lyn thus stands as a benchmark system, with profound implications for stellar evolution, black hole demographics, and strong-gravity astrophysics.
 II Results and Discussion
 II.1 Pulsation Properties and the Stellar Clock
-Analysis of the TESS photometry (Figure A1) reveals BE Lyn as a multi-periodic pulsator.
-We identify 37 significant frequencies, comprising 7 independent modes and 30 harmonics or combination frequencies (Figure A2, Table A1).
-The two dominant frequencies, $f_{0}=10.4309\pm 0.0005$ c/d and $f_{1}=13.4226\pm 0.0005$ c/d, exhibit a period ratio of $P_{1}/P_{0}=0.777$, unambiguously identifying them as the fundamental and first-overtone radial modes, respectively.
-The fundamental mode amplitude dominates the light curve, exceeding that of the first overtone by a factor of $\sim 45$ and the remaining independent modes by factors >700>700 (Figure A2).
-This clear amplitude hierarchy renders the $f_{0}$ pulsation a remarkably clean and stable clock, enabling the application of the classical OCO-C method to detect minute orbital perturbations with negligible contamination from multi-periodic effects.
-The five additional independent frequencies ($f_{2}$$f_{6}$), residing in the 9179-17 c/d range with low amplitudes, are likely non-radial p-modes.
+Analysis of the TESS photometry reveals BE Lyn as a multi-periodic pulsator.
+We identify 32 significant frequencies, comprising 7 independent modes and 25 harmonics or combination frequencies (see Appendix A for more details).
+The two dominant frequencies, $f_{0}=10.4309\pm 0.0005$ c/d and $f_{1}=13.4226\pm 0.0005$ c/d, exhibit a period ratio $P_{1}/P_{0}=0.777$, unambiguously identifying them as the fundamental and first-overtone radial modes, respectively (W.
+Xue et al., 2023).
+The amplitude of the fundamental mode dominates the light curve, exceeding that of the first overtone by a factor of $\sim 45$ and those of the remaining independent modes by factors >700>700.
+This clear amplitude hierarchy makes the $f_{0}$ pulsation an exceptionally clean and stable clock, enabling the classical OCO-C method to detect minute orbital perturbations with negligible contamination from multi-periodic effects.
+The five additional independent frequencies ($f_{2}$$f_{6}$), which lie in the 9179-17 c/d range with low amplitudes, are likely non-radial p-modes.
 This rich pulsation spectrum establishes BE Lyn as a promising target for future detailed asteroseismic modeling.
 II.2 Orbital Solution: An Extraordinarily Eccentric Binary
-The OCO-C analysis, incorporating a 39-year baseline of 445 times of maximum light, reveals a pronounced, long-period modulation superimposed on a secular period increase (Figure 1).
+The OCO-C analysis, incorporating a 39-year baseline of 445 times of maximum light, reveals a pronounced long-period modulation superimposed on a secular period increase (Figure 1).
 Global fitting with a model combining a quadratic term (intrinsic period variation) and a light-travel time effect (LTTE) term yields the orbital solution presented in Table 1.
@@
 Note. —
-The most striking result is the extraordinary orbital eccentricity of $e=0.9989^{+0.0008}_{-0.0021}$ (>0.9968>0.9968 at 95% confidence).
-With this value, BE Lyn claims the record for the highest reliably measured eccentricity in any binary system, surpassing previous benchmarks such as WDS 05354-3316 ($e=0.985\pm 0.002$ (A.
-Tokovinin et al., 2020)) and $\zeta$ Boo ($e=0.980450\pm 0.000064$ I.
-Waisberg et al.
-2024, 2025).
+The most striking result is the extraordinary orbital eccentricity: $e=0.9989^{+0.0008}_{-0.0021}$ (>0.9968>0.9968 at 95% confidence).
+With this value, BE Lyn sets a new record for the highest reliably measured eccentricity in any binary system, surpassing previous benchmarks such as WDS 05354-3316 ($e=0.985\pm 0.002$ (A.
+Tokovinin et al., 2020)) and $\zeta$ Boo ($e=0.980450\pm 0.000064$ (I.
+Waisberg et al., 2024, 2025)).
 The orbital period is $P_{\mathrm{orb}}=5811^{+29}_{-30}$ days ($\approx 15.9$ years).
-The associated mass function, $f(m)=0.005^{+0.046}_{-0.004}M_{\odot}$, and small projected semi-major axis ($a\sin i=1.1^{+1.3}_{-0.4}$ AU) are indicative of a massive, unseen companion.
-This discovery underscores the power of long-term, precise photometric monitoring of pulsating stars to unveil extreme binary architectures.
-II.3 Characterization of the Pulsating Primary
+The associated mass function, $f(m)=0.005^{+0.046}_{-0.004}M_{\odot}$, and small projected semi-major axis ($a\sin i=1.1^{+1.3}_{-0.4}$ AU) point to a massive, unseen companion.
+This discovery highlights the power of long-term, precise photometric monitoring of pulsating stars to uncover extreme binary architectures.
+II.3 Asteroseismic Properties of the Primary Star
 To place dynamical constraints in context, we determined the properties of the primary star through asteroseismic modeling.
@@
 Sun et al., 2023).
-The optimal model, selected by matching the observed $f_{0}$, $f_{1}$, and the intrinsic period variation rate $(1/P_{0})(dP_{0}/dt)$, is summarized in Table A2 and located on the Hertzsprung-Russell diagram in Figure 2.
-The best-fit model places the primary on the post-main sequence with a mass of $M_{1}=1.55M_{\odot}$, effective temperature $\log T_{\mathrm{eff}}=3.902$, and luminosity $\log(L/L_{\odot})=1.326$.
-The agreement between the observed $(1.02\pm 0.08)\times 10^{-8}$ yr-1 and modeled ($1.34\times 10^{-8}$ yr-1) period variation rates is noteworthy.
-Unlike some HADS in binaries that exhibit anomalously large period variations due to mass transfer (e.g., H.-F.
-Xue & J.-S.
-Niu, 2020; M.
-Lian et al., 2025), BE Lyn’s primary appears consistent with single-star evolution.
-This suggests the system is currently detached and has not undergone significant mass exchange.
+The optimal model, selected by matching the observed frequencies $f_{0}$ and $f_{1}$, is summarized in Table 3 and located on the Hertzsprung-Russell diagram in Figure 2.
 Figure 2: Hertzsprung–Russell diagram showing evolutionary tracks (colored solid lines from the zero-age main sequence, ZAMS, to post-MS) and the best-fit seismic model for BE Lyn’s primary (red star).
@@
 Colored dashed lines indicate tracks with a $0.01M_{\odot}$ mass step.
-II.4 A Nearby Black Hole in an Extreme Orbit
-Combining the orbital parameters with the primary mass ($M_{1}=1.55M_{\odot}$) introduces critical constraints from orbital stability: (i) the periastron separation must exceed the primary’s radius to avoid catastrophic collision; (ii) the primary’s Roche lobe radius must exceed the primary’s radius to avoid Roche-lobe overflow near the periastron.
-Applying these conditions, along with the mass function (Eq.
-A3), tightly constrains the system geometry (Figure 3).
-We derive an upper limit on the orbital inclination of $i\lesssim 4.0^{\circ}$, implying a lower limit on the companion mass of $M_{2}\gtrsim 17.5M_{\odot}$.
+The best-fit model places the primary on the post-main sequence, with a mass of $M_{1}=1.61M_{\odot}$, effective temperature $\log T_{\mathrm{eff}}=3.915$, and luminosity $\log(L/L_{\odot})=1.391$.
+Notably, the observed period variation rate, $(1.02\pm 0.08)\times 10^{-8}$ yr-1, differs from the model-predicted rate of $8.36\times 10^{-8}$ yr-1.
+This discrepancy may indicate ongoing mass transfer during the observed time interval.
+II.4 A Massive Compact Companion in an Extreme Orbit
+Combining the orbital parameters with the primary mass ($M_{1}=1.61M_{\odot}$) introduces a critical constraint from orbital stability: the periastron separation must exceed the primary’s radius to avoid catastrophic collision.
+This condition tightly constrains the system geometry (Figure 3).
+We derive an upper limit on the orbital inclination of $i\lesssim 10.1^{\circ}$, implying a lower limit on the companion mass of $M_{2}\gtrsim 2.5M_{\odot}$.
 Figure 3: Dynamical constraints on the unseen companion.
@@
 Bottom: Primary-companion separation at periastron versus $i$ (green curve).
-The intersection of the primary’s Roche lobe radius $R_{1,\mathrm{R}}$ (blue curve) and its physical radius $R_{1}$ (grey curve) sets the stringent upper limit $i\lesssim 4.0^{\circ}$ (dark blue dot), corresponding to the lower mass limit $M_{2}\gtrsim 17.5M_{\odot}$ (dark red dot).
+The intersection of the primary’s radius $R_{1}$ (grey curve) and primary-companion separation (green curve) sets the stringent upper limit $i\lesssim 10.1^{\circ}$ (dark green dot), corresponding to the lower mass limit $M_{2}\gtrsim 2.5M_{\odot}$ (dark red dot).
 The gray hatched region denotes parameter space excluded.
-A compact object of this mass cannot be supported by electron or neutron degeneracy pressure, ruling out a white dwarf or neutron star.
-The companion must therefore be a black hole.
-With a Gaia DR3 distance of approximately 250 pc (C.
+While natal kicks during an asymmetric supernova explosion can impart a relatively high eccentricity to binary systems (H.-T.
+Janka, 2013; S.
+Popov et al., 2025), the extreme eccentricity and semi-major axis of this system may require additional mechanisms to explain its later evolution.
+Interestingly, at periastron, the primary’s radius inevitably exceeds its Roche lobe radius (using the approximation of P.
+P.
+Eggleton (1983); see Figure 3), indicating that the primary must fill its Roche lobe, leading to mass transfer near periastron.
+[^1] Here, we thank P.
+Nagarajan et al.
+(2026) for pointing out an error in our previous calculation of the primary’s Roche lobe radius.
+This periastron mass transfer may have driven the elongation of the semi-major axis and the extreme eccentricity (A.
+Parkosidis et al., 2026, 2025), consistent with their unified semi-analytical framework and representing an extreme case of such evolution.
+A compact object with mass $M_{2}\gtrsim 2.5M_{\odot}$ cannot be supported by electron degeneracy pressure, ruling out a white dwarf.
+Its mass places it squarely spanning the mass gap between neutron stars and black holes.
+If the companion is a neutron star, it would challenge the current record for the most massive neutron star (PSR J0952-0607 at $2.35\pm 0.17M_{\odot}$ (R.
+W.
+Romani et al., 2022)).
+If instead it is a black hole, it would become the nearest known black hole to Earth, at a Gaia DR3 distance of approximately 250 pc (C.
 A.
 L.
-Bailer-Jones et al., 2021), this black hole becomes the nearest known to Earth, residing about half the distance of the previous record-holders Gaia BH1 (480 pc (K.
+Bailer-Jones et al., 2021)—roughly half the distance of the previous record-holders Gaia BH1 (480 pc (K.
 El-Badry et al., 2023)) and Gaia BH3 (590 pc (Gaia Collaboration et al., 2024)).
-Its proximity offers unparalleled opportunities for detailed follow-up study.
-The system’s most enigmatic feature—its extreme eccentricity—holds the key to its formation history.
-The absence of evidence for a dense stellar environment makes a dynamical scattering origin unlikely.
-Instead, the $e\approx 0.9989$ orbit is a natural signature of a strong natal kick imparted to the black hole during an asymmetric supernova explosion (H.-T.
-Janka, 2013; S.
-Popov et al., 2025).
-BE Lyn thus provides a pristine laboratory to constrain kick velocities and the physics of mass ejection in fallback supernovae, challenging models to account for the most eccentric black-hole binary known.
+Its proximity offers unparalleled opportunities for detailed follow-up observations, including astrometric monitoring and multi-wavelength searches for accretion signatures.
 II.5 Broader Implications and Future Prospects
-BE Lyn represents a system of singular importance across multiple astrophysical domains.
-First, it is the only known binary pairing a pulsating star with a quiescent black hole.
-This unique configuration opens the novel field of strong-gravity asteroseismology, where the pulsations of the primary could, in principle, probe the local spacetime metric perturbed by the black hole companion, especially during the brief, close periastron passages.
-Second, the discovery validates the pulsation-timing method as a powerful, independent technique for detecting dormant stellar-mass black holes in wide binaries.
-This approach is complementary to radial-velocity and astrometric surveys, being particularly sensitive to long-period systems where other methods require extended baselines.
-Key questions remain.
-Precise dynamical mass measurements via future astrometric missions (e.g., Gaia, Theia) or very-long-baseline interferometry could pin down $M_{2}$ and $i$.
+BE Lyn is a system of singular importance across multiple astrophysical frontiers.
+First, it is the only known binary comprising a pulsating star and a quiescent compact object spanning the mass gap.
+This unique configuration opens the door to strong-gravity asteroseismology: in principle, the primary’s pulsations could probe the local spacetime metric perturbed by the massive companion, especially during the brief, close periastron passages.
+Second, this discovery validates the pulsation-timing method as a powerful and independent technique for detecting dormant stellar-mass compact objects in wide binaries.
+Complementary to radial-velocity and astrometric surveys, this approach is particularly sensitive to long-period systems where other methods require extended baselines.
+Several key questions remain.
+Precise dynamical mass measurements via future astrometric missions (e.g., Gaia, Theia) or very-long-baseline interferometry could determine $M_{2}$ and $i$ unambiguously.
 Multi-wavelength searches for accretion signatures or bow-shock emission are warranted, despite the wide separation.
-Theoretical work is urgently needed to simulate binary evolution pathways that yield such extreme eccentricities post-supernova.
+Theoretical work is urgently needed to model binary evolution pathways that produce such extreme eccentricities post-supernova.
 In conclusion, BE Lyn is not merely another binary system; it is a benchmark object.
-It challenges our understanding of black hole formation kicks, offers a new tool for black hole discovery, and provides a unique cosmic laboratory at our galactic doorstep.
+It challenges our understanding of black hole formation kicks, offers a new tool for discovering compact objects, and provides a unique cosmic laboratory at our galactic doorstep.
 III Conclusion
 We have established BE Lyncis as an extraordinary binary system.
-The primary is a post-main sequence high-amplitude $\delta$ Scuti star with a mass of $1.55\,M_{\odot}$.
-Its unseen companion orbits with an unprecedented eccentricity of $e=0.9989^{+0.0008}_{-0.0021}$ (95% confidence >>0.9968), making this the most eccentric binary known to date.
-This extreme orbit, coupled with dynamical constraints that demand a companion mass $M_{2}\gtrsim 17.5\,M_{\odot}$ at an orbital inclination $i\lesssim 4.0^{\circ}$, unequivocally identifies the companion as a stellar-mass black hole.
-With a distance of approximately 250 pc, this black hole is the closest known to Earth.
+The primary is a post-main-sequence, high-amplitude $\delta$ Scuti star with a mass of $1.61\,M_{\odot}$.
+Its unseen companion follows an orbit with an eccentricity of $e=0.9989^{+0.0008}_{-0.0021}$ (95% confidence >0.9968>0.9968), the most extreme reliably measured in any binary system to date.
+Dynamical constraints, rooted in orbital stability, limit the inclination to $i\lesssim 10.1^{\circ}$ and demand a companion mass $M_{2}\gtrsim 2.5\,M_{\odot}$, placing it squarely spanning the mass gap between neutron stars and black holes.
 The discovery of BE Lyn carries profound implications across several astrophysical frontiers.
-First, its record-breaking eccentricity serves as a direct observational constraint on the strength of natal kicks imparted during asymmetric supernova explosions, challenging models of black hole formation.
-Second, it is the first known binary comprising a pulsating star and a quiescent black hole, creating a unique laboratory for strong-gravity asteroseismology.
-Third, this work validates the pulsation-timing method—using the precise clocks of $\delta$ Scuti stars—as a novel, independent pathway to discover dormant black holes in wide binaries, thereby opening a new observational channel in the quest to unveil the hidden compact-object population.
-Key questions regarding the exact black hole mass, the detailed evolutionary pathway leading to such an extreme configuration, and the system’s ultimate fate remain.
-Future multi-wavelength observations, astrometric monitoring, and theoretical modeling of supernova kicks in binary systems will be essential to fully understand the origins and implications of this remarkable cosmic benchmark.
-Materials & Correspondence
-Correspondence and requests for materials should be addressed to Jia-Shu Niu. (email: jsniu@sxu.edu.cn).
+First, its record-breaking eccentricity provides a stringent observational benchmark for models of natal kicks in asymmetric supernovae, challenging our understanding of compact-object formation and evolution.
+Second, it represents the first known binary system to pair a pulsating star with a quiescent compact object spanning the mass gap, opening the door to strong-gravity asteroseismology—a novel probe of spacetime around massive companions.
+Third, our work validates the pulsation-timing method, which harnesses the precise clocks of $\delta$ Scuti stars, as an independent and powerful pathway to discover dormant compact objects in wide binaries.
+This technique complements existing radial-velocity and astrometric surveys and offers a new observational window into the hidden population of neutron stars and black holes.
+Fundamental questions remain.
+The exact mass and nature of the companion—whether the most massive neutron star known or the nearest black hole to Earth—await definitive measurement.
+The evolutionary pathway that produced such an extreme configuration, and the ultimate fate of the system, are yet to be understood.
+Future multi-wavelength observations, astrometric monitoring (e.g., with Gaia and future missions), and theoretical modeling of supernova kicks and binary evolution will be essential to fully unravel the origins and implications of this remarkable cosmic benchmark.
+Acknowledgments
 H.F.X.
@@
 All the TESS data used in this paper can be found in MAST: http://dx.doi.org/10.17909/t9-nmc8-f686 (catalog 10.17909/t9-nmc8-f686).
-{contribution}
-J.S.N.
-and H.F.X.
-designed the research plan, performed the calculation, and wrote the manuscript.
-Y.Z.
-reviewed relevant literature, collected the TML, performed the frequency analysis, and wrote the initial version of Sec.
-1 and part of Sec.
-2.
-Competing Interests
-The authors declare that they have no competing financial interests.
-Appendix A Methods
-A.1 Photometric Data and Frequency Analysis
+Appendix A Photometric Data and Frequency Analysis
 BE Lyn was observed by the Transiting Exoplanet Survey Satellite (TESS) in Sector 21 (January 2020) with a 120-second cadence.
@@
 G.
-Stassun et al., 2019).
+Stassun et al., 2019) (see in Figure 4).
+Figure 4: Light curve of BE Lyn from TESS Sector 21. The lower panel shows a one-day zoom for clarity.
 We extracted pulsation frequencies using an iterative prewhitening procedure (J.-S.
@@
 where $m_{0}$ is the mean magnitude, and $a_{i}$, $f_{i}$, and $\phi_{i}$ are the amplitude, frequency, and phase of each component.
-Frequencies were accepted as significant if their signal-to-noise ratio (S/N) exceeded 5.6 (J.-S.
-Niu & H.-F.
-Xue, 2022; J.-S.
-Niu et al., 2023; J.-S.
-Niu & H.-F.
-Xue, 2024).
-After removing aliases, we identified 37 significant peaks, consisting of 7 independent modes and 30 harmonics or linear combinations (Table A1).
-A.2 Compilation and Processing of Times of Maximum Light
+Frequencies were accepted as significant if their signal-to-noise ratio (S/N) exceeded 5.6 (W.
+Zong et al., 2016, 2018).
+After removing aliases, we identified 32 significant peaks, consisting of 7 independent modes and 25 harmonics or linear combinations (see in Figure 5 and Table 2).
+Figure 5: Frequency spectrum of BE Lyn.
+The main panel (0100c/d100\ $\mathrm${c/d}) marks the fundamental ($f_{0}$, red) and first-overtone ($f_{1}$, green) modes.
+The inset highlights the low-amplitude, non-radial mode region ($\sim 9-17\ \mathrm{c/d}$) with a linear y-axis.
+Table 2: Significant Frequencies Detected in BE Lyn from TESS Sector 21.
+NO. f(c/d)f\ ($\mathrm${c/d}) $\sigma_{f}\ (\mathrm{c/d})$ a(mmag)a\ ($\mathrm${mmag}) $\sigma_{a}\ (\mathrm{mmag})$ S/N Marks
+F1 10.4309 0.0005 116 3 42.4 $f_{0}$
+F2 20.8618 0.0005 38.9 0.9 42.1 $2f_{0}$
+F3 31.2926 0.0005 14.3 0.3 42.4 $3f_{0}$
+F4 41.7236 0.0005 6.0 0.1 44.2 $4f_{0}$
+F5 13.4226 0.0005 2.59 0.07 37.8 $f_{1}$
+F6 52.1544 0.0005 2.57 0.06 40.3 $5f_{0}$
+F7 62.5852 0.0005 1.41 0.03 41.3 $6f_{0}$
+F8 23.8534 0.0005 1.04 0.03 38.4 $f_{0}+f_{1}$
+F9 73.0156 0.0006 0.79 0.02 34.6 $7f_{0}$
+F10 2.9914 0.0007 0.71 0.03 27.4 $f_{0}+f_{1}$
+F11 34.2846 0.0008 0.50 0.02 27.0 $2f_{0}+f_{1}$
+F12 83.4470 0.0006 0.47 0.01 31.8 $8f_{0}$
+F13 7.440 0.001 0.37 0.02 21.3 $2f_{0}-f_{1}$
+F14 44.7150 0.0008 0.29 0.01 24.4 $3f_{0}+f_{1}$
+F15 93.8765 0.0009 0.26 0.01 23.8 $9f_{0}$
+F16 55.147 0.001 0.16 0.01 17.7 $4f_{0}+f_{1}$
+F17 9.701 0.003 0.15 0.02 7.2 $f_{2}$
+F18 104.309 0.002 0.14 0.01 14.0 $10f_{0}$
+F19 17.869 0.002 0.11 0.01 10.5 $3f_{0}-f_{1}$
+F20 65.579 0.002 0.10 0.01 10.4 $5f_{0}+f_{1}$
+F21 10.684 0.003 0.10 0.02 6.0 $f_{3}$
+F22 9.377 0.003 0.09 0.01 6.0 $f_{4}$
+F23 114.740 0.002 0.076 0.008 9.8 $11f_{0}$
+F24 14.009 0.003 0.08 0.01 6.4 $f_{5}$
+F25 19.810 0.003 0.08 0.01 6.2 $f_{0}+f_{4}$
+F26 20.133 0.003 0.07 0.01 5.9 $f_{0}+f_{2}$
+F27 21.112 0.004 0.07 0.01 5.7 $f_{0}+2f_{3}$
+F28 76.006 0.003 0.067 0.009 7.5 $6f_{0}+f_{1}$
+F29 16.336 0.003 0.07 0.01 6.8 $f_{6}$
+F30 28.299 0.003 0.058 0.009 6.1 $4f_{0}-f_{1}$
+F31 38.729 0.003 0.052 0.008 6.1 $5f_{0}-f_{1}$
+F32 86.438 0.004 0.045 0.008 5.8 $7f_{0}+f_{1}$
+Appendix B Compilation and Processing of Times of Maximum Light
 The light curve is dominated by the fundamental mode ($f_{0}$), with the first overtone ($f_{1}$) contributing less than 3% of its amplitude (W.
@@
 (2011) published without uncertainties, we adopted a conservative error of 0.0001 d, consistent with their reported precision and twice the mean uncertainty of our TESS-based measurements.
-A.3 Orbital Modeling via MCMC
-The OCO-C residuals were modeled with a function combining a quadratic term (intrinsic period variation) and a light-travel time effect (LTTE) term for orbital motion (H.-F. Xue & J.-S. Niu, 2020):
-$C=\mathrm{BJD}_{0}+P_{0}E+\frac{1}{2}\beta E^{2}+A\left[\sqrt{1-e^{2}}\sin{\phi}\cos{\omega}+\cos{\phi}\sin{\omega}\right],$ (A2)
+Appendix C Orbital Modeling via MCMC
+The OCO-C residuals were modeled with a function combining a quadratic term (intrinsic period variation) and a light-travel time effect (LTTE) term for orbital motion (M.
+Paparo et al., 1988; H.-F.
+Xue & J.-S.
+Niu, 2020):
+[…]
 where $\phi$ is the eccentric anomaly from Kepler’s equation, $\mathrm{BJD}_{0}$ is the reference epoch, $P_{0}$ the pulsation period, $\beta$ the linear period variation rate, $A\equiv a\sin i/c$ the projected semi-major axis, $e$ the orbital eccentricity, $\omega$ the argument of periastron, $P_{\mathrm{orb}}$ the orbital period, and $T_{0}$ the time of periastron passage.
@@
 The best-fit model ($\chi^{2}/\mathrm{d.o.f.}=35.23$) is shown in Figure 1, with parameters listed in Table 1.
-A.4 Asteroseismic Modeling of the Primary Star
+Appendix D Asteroseismic Modeling of the Primary Star
 We estimated the primary star’s properties by matching the observed pulsation characteristics to stellar models.
@@
 Xue (2025).
-The optimal model was selected by minimizing $\chi^{2}$ across three observables: the fundamental frequency ($f_{0}$), first-overtone frequency ($f_{1}$), and the intrinsic period variation rate $(1/P_{0})(dP_{0}/dt)$.
+The optimal model was selected by minimizing $\chi^{2}$ across two observables: the fundamental frequency ($f_{0}$) and first-overtone frequency ($f_{1}$).
 This methodology follows established asteroseismic analyses of HADS (J.-S.
@@
 Xue et al., 2018, 2022).
-The best-fit model parameters are given in Table A2.
-A.5 Dynamical Constraints on the Companion
+The best-fit model parameters are given in Table 3.
+Table 3: Parameters of the best-fit seismic model for BE Lyn’s primary star.
+Parameter Best-fit Model Observed
+$f_{0}$ (c/d) 10.4310 $10.4309\pm 0.0005$
+$f_{1}$ (c/d) 13.4226 $13.4226\pm 0.0005$
+$(1/P_{0})(dP_{0}/dt)$ (yr-1) $8.36\times 10^{-8}$ $(1.02^{+0.08}_{-0.08})\times 10^{-8}$
+$\log T_{\mathrm{eff}}$ 3.915
+$\log(L/L_{\odot})$ 1.391
+$\log g$ 3.870
+$\log(R/R_{\odot})$ 0.387
+Age (Gyr) 1.44
+Mass ($M_{\odot}$) 1.61
+Appendix E Dynamical Constraints on the Companion
 The companion’s nature was inferred from the mass function,
-$f(M_{2})=\frac{(M_{2}\sin i)^{3}}{(M_{1}+M_{2})^{2}}=\frac{4\pi^{2}}{GP_{\mathrm{orb}}^{2}}(a\sin i)^{3},$ (A3)
-combined with the constraints that the periastron separation $d_{\mathrm{peri}}=a(1-e)(M_{1}+M_{2})/M_{2}$ must exceed the primary’s radius $R_{1}$, and the Roche lobe radius $R_{1,\mathrm{R}}$ must exceed the primary’s radius $R_{1}$.
-The Roche lobe radius was calculated using the approximation of P.
-P.
-Eggleton (1983).
-This yields an upper limit on the inclination ($i\lesssim 4.0^{\circ}$) and a corresponding lower limit on the companion mass ($M_{2}\gtrsim 17.5M_{\odot}$), as shown in Figure 3.
-Appendix B Additional Figures
-Figure A1: Light curve of BE Lyn from TESS Sector 21. The lower panel shows a one-day zoom for clarity.
-Figure A2: Frequency spectrum of BE Lyn.
-The main panel (0100c/d100\ $\mathrm${c/d}) marks the fundamental ($f_{0}$, red) and first-overtone ($f_{1}$, green) modes.
-The inset highlights the low-amplitude, non-radial mode region ($\sim 9-17\ \mathrm{c/d}$) with a linear y-axis.
-Appendix C Additional Tables
-Table A1: Frequency solution for BE Lyn from TESS Sector 21.
-NO. f(c/d)f\ ($\mathrm${c/d}) $\sigma_{f}\ (\mathrm{c/d})$ a(mmag)a\ ($\mathrm${mmag}) $\sigma_{a}\ (\mathrm{mmag})$ S/N Marks
-F1 10.4309 0.0005 116 3 42.4 $f_{0}$
-F2 20.8618 0.0005 38.9 0.9 42.1 $2f_{0}$
-F3 31.2926 0.0005 14.3 0.3 42.4 $3f_{0}$
-F4 41.7236 0.0005 6.0 0.1 44.2 $4f_{0}$
-F5 13.4226 0.0005 2.59 0.07 37.8 $f_{1}$
-F6 52.1544 0.0005 2.57 0.06 40.3 $5f_{0}$
-F7 62.5852 0.0005 1.41 0.03 41.3 $6f_{0}$
-F8 23.8534 0.0005 1.04 0.03 38.4 $f_{0}+f_{1}$
-F9 73.0156 0.0006 0.79 0.02 34.6 $7f_{0}$
-F10 2.9914 0.0007 0.71 0.03 27.4 $f_{0}+f_{1}$
-F11 34.2846 0.0008 0.50 0.02 27.0 $2f_{0}+f_{1}$
-F13 83.4470 0.0006 0.47 0.01 31.8 $8f_{0}$
-F14 7.440 0.001 0.37 0.02 21.3 $2f_{0}-f_{1}$
-F15 44.7150 0.0008 0.29 0.01 24.4 $3f_{0}+f_{1}$
-F16 93.8765 0.0009 0.26 0.01 23.8 $9f_{0}$
-F19 55.147 0.001 0.16 0.01 17.7 $4f_{0}+f_{1}$
-F20 9.701 0.003 0.15 0.02 7.2 $f_{2}$
-F22 104.309 0.002 0.14 0.01 14.0 $10f_{0}$
-F24 17.869 0.002 0.11 0.01 10.5 $3f_{0}-f_{1}$
-F25 65.579 0.002 0.10 0.01 10.4 $5f_{0}+f_{1}$
-F26 10.684 0.003 0.10 0.02 6.0 $f_{3}$
-F27 9.377 0.003 0.09 0.01 6.0 $f_{4}$
-F28 114.740 0.002 0.076 0.008 9.8 $11f_{0}$
-F29 14.009 0.003 0.08 0.01 6.4 $f_{5}$
-F30 19.810 0.003 0.08 0.01 6.2 $f_{0}+f_{4}$
-F31 20.133 0.003 0.07 0.01 5.9 $f_{0}+f_{2}$
-F32 21.112 0.004 0.070 0.01 5.7 $f_{0}+2f_{3}$
-F33 76.006 0.003 0.067 0.009 7.5 $6f_{0}+f_{1}$
-F34 16.336 0.003 0.07 0.01 6.8 $f_{6}$
-F35 28.299 0.003 0.058 0.009 6.1 $4f_{0}-f_{1}$
-F36 38.729 0.003 0.052 0.008 6.1 $5f_{0}-f_{1}$
-F37 86.438 0.004 0.045 0.008 5.8 $7f_{0}+f_{1}$
-Table A2: Parameters of the best-fit seismic model for BE Lyn’s primary star.
-Parameter Best-fit Model Observed
-$f_{0}$ (c/d) 10.4288 $10.4309\pm 0.0005$
-$f_{1}$ (c/d) 13.4243 $13.4226\pm 0.0005$
-$(1/P_{0})(dP_{0}/dt)$ (yr-1) $1.34\times 10^{-8}$ $(1.02^{+0.08}_{-0.08})\times 10^{-8}$
-$\log T_{\mathrm{eff}}$ 3.902
-$\log(L/L_{\odot})$ 1.326
-$\log g$ 3.865
-$\log(R/R_{\odot})$ 0.382
-Age (Gyr) 1.59
-Mass ($M_{\odot}$) 1.55
+$f(M_{2})=\frac{(M_{2}\sin i)^{3}}{(M_{1}+M_{2})^{2}}=\frac{4\pi^{2}}{GP_{\mathrm{orb}}^{2}}(a\sin i)^{3},$ (E1)
+combined with the constraints that the periastron separation $d_{\mathrm{peri}}=a(1-e)(M_{1}+M_{2})/M_{2}$ must exceed the primary’s radius $R_{1}$.
+This yields an upper limit on the inclination ($i\lesssim 10.1^{\circ}$) and a corresponding lower limit on the companion mass ($M_{2}\gtrsim 2.5M_{\odot}$), as shown in Figure 3.
+Appendix F Additional Tables
+All the TML collected and used in this work are listed in Table 4.
 […]
-Table A3: Times of maximum light of BE Lyn.
+Table 4: Times of maximum light of BE Lyn.
 The numbers in the fourth, eighth, and twelfth columns indicate the different data sources: (1) T.